OpenStudy (anonymous):

When we are talking about the imaginary solutions to a polynomial, can we use the word "zeros" or do we only use the word solutions since the imaginary answers to not touch the x-axis?

OpenStudy (agent0smith):

Imaginary zeroes.

OpenStudy (amistre64):

\[(\sqrt-2)^2+2 = 0\] \[-2+2 = 0\] they are still zeros of the function

OpenStudy (anonymous):

so just as you refer to solutions of a quadratic equations as zeros, solutions, answers, roots: you can use the same nomenclature for the imaginary solutions of a polynomial ?

OpenStudy (agent0smith):

Well a quadratic is a polynomial too, and it can have imaginary zeroes.

OpenStudy (anonymous):

Good point. Thanks I'll close the question and ask a new one!

OpenStudy (amistre64):

the geometric interpretation leads to the x intercepts being defined as "zeros" ; but a polynomial is not restricted to a geometric interpretaion